OpenStudy (anonymous):

i'm trying to solve the x-intercept and having a hard time with it 5x+8y=40

OpenStudy (anonymous):

an x-intercept is a point in the equation where the y-value is zero, and a y-intercept is a point in the equation where the x-value is zero. Using these definitions y=0 for the x intercept. So, 5x+8(0) = 40 5x+0=40 5x=40 x=40/5 x=8 - Answer The x intercept is 8

OpenStudy (anonymous):

thank you so much somehow i was coming up with a mixed number it was a disaster